The volume of a sphere is V=4/3(pi)r^3. Find the radius to the nearest hundred of a sphere with a volume of 15 inches^3
Answer by avani(25)

The volume of a sphere is V=4/3(pi)r^3
V=4/3πr^3
15=4/3πr^3
((15)/(4/3π))^1/3=r
1.53=r